Sarah is 5 years younger than 3 times Joes age, the sum of their ages is 43. What are both of their ages?
Answer by Maths68(1474)

If we subtract 5 from three times of joes age, Joe and Sarah will be of same age
s=3j-5…………(1)
Sum of their ages is 43
s+j=43
s=43-j…………….(2)
Put the value of s from (2) to (1)
s=3j-5
43-j =3j-5
43-j+5=3j
48-j=3j
48=3j+j
4j=48
Divide by 4 both sides
4j/4=48/4
j=12
Put the value of j in (2), we have
s=43-12
s=31
Sarah is 31 and Joe is 12 years old